SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: February 27, 2008
Shift Taker: Graham Jackson
-----------
General Features:
(theta,phi)

Warm crystals in hitmap	(52,63), (56,63), (51,64)
Warm crystals in 0.01 GeV hitmap (17,43), (30,45), (32,46), (36,64)
Warm crystals in 0.1 GeV hitmap (30,45), (32,46), (36,64)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.1 GeV hitmap (29,84) 
Cool crystal  in 1 GeV hitmap    		

[1] =  TC Damage:OOS/OOS ~10^-3 in 0-5 all fibers. (And usually FLinkA/B/C Timeout is high too).

0079976 [1] and also TC Damage:OOS/OOS ~10^-3 in 3-6 all fibers. 
0079977 [1]
0079978 [1]
0079979 [1]
0079980 [1]
0079981 ok.
0079982 [1]
0079983 [1]. Evidence of a noisy crystal on digi timing plot at LogE ~ 1.
0079984 Short cosmics run.  No damage.
0079991 Short cosmics run.  No damage.
0079992 Short cosmics run.  No damage.  (Probably, since there were only 4 pages in the file, and only 2 in the backup.)
0079993 Cosmics run.  No Damage.
0079994 Short cosmics run.  No damage.
0079995 Evidence of noisy channels to left and right of digi timing peak, at logE~ -0.5 and 1 respectively.
 	  	   And the energy scale seems odd on the 1GeV hit map!?...Too warm?


Crystal Spectra:


Lightpulser:
4 channels dead, 1 channels out of time. Updated on: 27-February 2008 16:04 PST


 

  Links

  • Shift Report Index
  • Back to the Shifter instructions